# DDaTR: Dynamic Difference-aware Temporal Residual Network for Longitudinal Radiology Report Generation

### v1

@@ Abstract
Radiology Report Generation (RRG) automates the creation of radiology reports from medical imaging, enhancing the efficiency of the reporting process. Longitudinal Radiology Report Generation (LRRG) extends RRG by incorporating the ability to compare current and prior exams, facilitating the tracking of temporal changes in clinical findings. Existing LRRG approaches only extract features from prior and current images using a visual pre-trained encoder, which are then concatenated to generate the final report. However, these methods struggle to effectively capture both spatial and temporal correlations during the feature extraction process. Consequently, the extracted features inadequately capture the information of difference across exams and thus underrepresent the expected progressions, leading to sub-optimal performance in LRRG. To address this, we develop a novel dynamic difference-aware temporal residual network (DDaTR). In DDaTR, we introduce two modules at each stage of the visual encoder to capture multi-level spatial correlations. The Dynamic Feature Alignment Module (DFAM) is designed to align prior features across modalities for the integrity of prior clinical information. Prompted by the enriched prior features, the dynamic difference-aware module (DDAM) captures favorable difference information by identifying relationships across exams. Furthermore, our DDaTR employs the dynamic residual network to unidirectionally transmit longitudinal information, effectively modelling temporal correlations. Extensive experiments demonstrated superior performance over existing methods on three benchmarks, proving its efficacy in both RRG and LRRG tasks. Our code is published at https://github.com/xmed-lab/DDaTR
Index Terms:
@@ I Introduction
  • We propose a novel dynamic difference-aware temporal residual network (DDaTR) to better encode longitudinal information by adaptively learning the multi-level abstractions of different prior features and modelling the unidirectional information flow.
@@
II Related Work
II-A Radiology Report Generation
@@
II-B Longitudinal Radiology Report Generation
@@
Although adding longitudinal features could alleviate the hallucinations about unobserved priors when training language models for report generation, there are still limitations in capturing the important longitudinal correlations. Firstly, they overlook the different contributions of previous image content to current reports across different samples in real clinical scenarios (e.g., different viewpoints, limited longitudinal information or varying time intervals between examinations). Most of them only considered the frontal view scans between previous and current periods. This limitation leads to insufficient global temporal correlations. Secondly, previous feature fusion is used after visual encoder. Specifically, they use general visual pre-training encoders (e.g.,ResNet [30], SwinTransformer [31]) to extract visual features of each images independently, and then they leverage cross-attention based modules or transformer blocks to encode longitudinal features. However, the process of encoding image features does not consider the relationships between multi-level features of longitudinal images, especially the low-level features including important details and texture information. This limitation leads to inaccurate local temporal correlations. Thirdly, previous methods fail to account for the prior knowledge in previous report during the visual feature extraction. This report documents the patient’s clinical findings and diagnoses, reflecting the underlying clinical intent of the examination, which is important for perceiving prior visual semantics.
III Proposed Method
@@
III-A Problem Definition and Our Baseline Framework
@@
$\small F_{img}=E_{img}\left(I_{n}\right)\vspace{-2mm}$ (1)
@@
$F_{img}=\begin{cases}E_{ddatr}\left(I_{n}\right)&\text{if }n=1\\ E_{ddatr}\left(I_{n},I_{n-1},R_{n-1}\right),&\text{if }n>1\end{cases}\vspace{-1mm}$ (4)
@@
III-B Overview of Our Proposed DDaTR
@@
where $\hat{F}_{prior}^{m-1}$ represents the final output feature from the $m-1$ stage and $E_{prior}^{m}$ denotes the $m^{th}$ stage of prior visual encoder. For each stage of the prior visual encoder, we introduce the Dynamic Feature Alignment Module (DFAM) to dynamically fuse the image feature $F_{prior}^{m}$ and text feature $F_{txt}$, resulting in the aligned output feature $F_{align}^{m}$.
@@
where the mechanics of the $DFAM(\cdot)$ are detailed in Sec. III-C. Next, this aligned output is added to $F_{prior}^{m}$ via a residual connection to obtain the final prior visual feature. The formula is given by:
@@
The mechanics of the $DDAM(\cdot)$ are detailed in Sec. III-D. This output is also added to $F_{cur}^{m}$ using a residual connection, subsequently feeding into the next stage. Ultimately, this process culminates in the acquisition of the current visual feature $\hat{F}_{cur}^{m}$. The equation can be expressed as:
@@
TABLE I: Comparison with exisiting report generation methods on MIMIC-CXR, Longitudinal-MIMIC and IU-XRay. * indicates results evaluated by us. For the MIMIC-CXR and IU-Xray, we adopt the evaluation outlined by RECAP[8], using the Macro score for CE metrics. For the Longitudinal-MIMIC, we follow L-R2Gen [3], calculating the Micro score for CE metrics.
@@
III-C Dynamic Feature Alignment Module
@@
III-D Dynamic Difference-aware Module
@@
Secondly, the dynamic fusion part of the DDAM is the same as DFAM, designed to ensure that the difference-aware features do not overwhelm the current visual features while allowing a controlled flow of longitudinal information into the next stage. This mudule dynamically integrates prior information, allowing the model to adaptively select the most relevant features. Specifically, $F_{cur}^{m}$ represents the vasual features, $F_{da}$ means the difference-aware visual features including longitudinal features. When we fuse the $F_{cur}$ and $F_{da}$, we dynamically add $F_{da}$ to $F_{cur}$ using a gating mechanism. $f_{dy}$ represents our gating module, and the fusion output is $F_{dy}=f_{dy}(F_{da})$, where $F_{dy}$ represents the output weights generated by our gating module. Then, using ‘$\alpha$’ indicates whether the input has a prior examination; it is set to 1 if a prior exists, and 0 otherwise. So the fused process is as follows:
@@
IV Experiments
IV-A Datasets and Metrics
@@
TABLE II: Performance Comparison on three different visual backbones: ResNet101, SwinT-B, and EfficientNetv2. Vanilla represents that visual features were extracted using original pretrained models, coupled with the frozen pretrained BERT for text feature extraction, with subsequent concatenation of these features.
Datasets Visual Backbone Methods Clinical Efficacy Natural Language Generation
@@
MIMIC -CXR ResNet101 Vanilla 0.395 0.426 0.389 0.108 0.156 0.267
@@
Longitudinal -MIMIC ResNet101 Vanilla 0.487 0.544 0.486 0.101 0.151 0.261
@@
TABLE III: Ablation Study. w/o MSF indicates without multi-stage fusion. w/o DAM indicates without difference-aware module. w/o FAM indicates without feature aligment module. w/o DF indicates without dynamic fusion of both DFAM and DDAM.
@@
Experiment Settings. Our code, based on the Pytorch framework [47], was trained using a single NVIDIA RTX A40 GPU. For all methods used in our comparison, we strictly followed the default configurations reported in their respective original papers without any modifications. To ensure fair and consistent evaluation, we adopted the same training settings as the baseline PromptMRG across all our experiments, including those conducted on different datasets and with various visual backbones. Specifically, all input images were resized to 256 × 256, followed by data augmentations such as 224 × 224 random cropping and random rotation within ±5 degrees. The maximum input text length was set to 100 words. We employed the AdamW optimizer [48] with a weight decay of 0.05. The initial learning rate was set to 5e-5, and all models were trained for 10 epochs with a batch size of 16.
IV-B Experimental Results
Comparison on both RRG (MIMIC-CXR, IU-XRay) and LRRG (Longitudinal-MIMIC) benchmarks. We compare our method with a wide range of classical and SOTA methods on MIMIC-CXR, IU-XRay and Longitudinal-MIMIC, and report their results in Tab. I. For the MIMIC-CXR dataset, we follow the evaluation methodology established by RECAP, calculating the Macro score for the CE metrics. For the IU-Xray dataset, we use the same setting as PromptMRG that did not conduct any training or fine-tuning on it, but directly tested our model which trained on MIMIC-CXR. For the Longitudinal-MIMIC dataset, our approach follows the evaluation methodology used in LR2Gen, where the Micro score is calculated for the CE metrics. By comparison, our method achieves state-of-the-art (SOTA) performance on CE metrics, significantly outperforming other approaches on both RRG and LRRG. We performed a statistical significance analysis on the MIMIC-CXR test set by calculating the p-value of the F1 scores across all longitudinal cases. The result is $P_{F1}$=0.0415 <0.05 comparing with previous SOTA promptmrg, indicating that the performance improvement of our method is statistically significant. This not only demonstrates that utilizing longitudinal information for report generation tasks yields reports that more closely align with the groundtruth, but also highlights our method’s capability to more accurately capture the diagnostic intent of the radiologists, resulting in more precise report diagnoses. For example, our method with SwinT-B on MIMIC-CXR achieves 0.441 of F1_macro CE score with an absolute improvement of 0.048. Our method with SwinT-B on Longitudinal-MIMIC achieves 0.527 of F1_micro CE score with an absolute improvement of 0.035.
TABLE IV: Performance Metrics on MIMIC-CXR Dataset.Comparison with Large language models.
@@ IV-B Experimental Results
TABLE V: Performance on ReXrank Dataset. Comparison with existing specialist RRG Models.
@@
Figure 7: (a) Illustrates different test settings. Each column represents a specific test setting, with nindicating the proportion of longitudinal reports used. The total number of samples for each setting is shown at the top of each column. (b) Performance Comparison on different test settings.
@@
We analyse this as follows: a review of existing methods reveals that higher NLG scores often result from enhancements in the decoder. Thus, under similar data processing and training strategies, the NLG scores largely depends on the ability of the model’s decoder, which includes aspects such as the architecture, the tokenizer, and the pre-trained models. Our research primarily focuses on how to encode longitudinal information effectively, aligning input features more closely with medical intent. Consequently, we did not make specific structural modifications to the decoder but instead followed the decoder used in PromptMRG. Furthermore, our method also outperforms previous SOTA on IU X-Ray, which does not contain longitudinal information. This further demonstrates that our model not only substantially improve longitudinal report generation, but also maintain the advanced performance in single-report scenarios. Moreover, the strong zero-shot results on an external dataset highlight the robustness and generalizability of our approach.
@@
Comparison on different visual backbones. Our proposed algorithmic framework is adaptable to all multi-scale visual feature extraction networks. To validate its effectiveness, we conducted comparative experiments using three prominent visual backbones: ResNet101 [30], SwinT-B [31], and EfficientNetv2 [59]. These networks were selected because they are commonly employed for visual feature extraction in prior RRG research. We compared the performance outcomes between using standard vanilla visual networks and our method under the same baseline framework. In the vanilla approach, visual features were extracted using the original network architectures (e.g., ResNet101) and also fine-tuned during training using the same strategy as ours. For textual features, a pretrained BERT model was also used in a frozen state, and the visual and textual features were subsequently concatenated for downstream tasks. Experimental results, presented in Tab. II, demonstrate that our method outperforms the vanilla encoder across all three backbones by a significant margin.
@@
Comparison on ReXrank. To compare our method with the latest approaches in RRG, we evaluated DDaTR on the official test set of MIMIC-CXR provided by the ReXrank leaderboard [60], which including the comprehensive metrics, such as BERTScore[61], SembScore[32], and GREEN[62]. The results are presented in Tab. V. Our model ranks second on the leaderboard. While MedVersa [58] significantly outperforms all other methods on the leaderboard, it is trained on 10 public datasets and built upon LLM. Due to the substantial differences in both data amount and model size, direct comparison is not entirely fair. Compared to other task-specific LLMs that are explicitly trained on RRG data—such as LLM-CXR[55], CheXagent[57], RaDialog[54] and MAIRA-2[56]—our method achieves consistently superior performance across all evaluation metrics, further demonstrating the effectiveness of our approach under a more lightweight setting.
@@
Comparison on complexity. To compare model complexity, we report the number of trainable parameters for our method and the baseline PromptMRG as follows: PromptMRG (220M), Ours with Swin-T (360M), Ours with EfficientNetV2 (247M), and Ours with ResNet101 (419M). Notably, when using EfficientNetV2, our model maintains a comparable parameter count (247M) to PromptMRG (220M), yet still achieves significantly better performance as shown in Tab. II. This indicates that the performance gains of our method are not merely attributed to increased model complexity. In fact, in Tab. II Longitudinal-MIMIC dataset, our network with ResNet101 has the highest number of parameters but does not perform as well as the other backbones, further emphasizing that simply increasing model complexity does not guarantee better longitudinal modeling performance. Furthermore, PromptMRG is designed for single-period input, while our method handles longitudinal data, which is acceptable to introduce some additional parameters.
@@
Comparison on different proportions of longitudinal reports. We conducted assessments not only on standard MIMIC-CXR benchmark but also under various settings with different proportions of longitudinal reports. As shown in Fig. 7 (a), we split the testset of MIMIC-CXR. 0% indicates that all test data consists exclusively of reports from patients’ initial visits without longitudinal information. Conversely, 100% means that the test set includes all reports. At 50%, the test set is composed by randomly selecting half of the reports with prior exams. In the above three testing settings, we evaluated the performance of different encoding methods within the baseline frameworks. Specifically, we tested the effectiveness of three prominent visual backbones—ResNet101, SwinT-B, and EfficientNetv2—using both the standard vanilla approach and our method, alongside another pretrained encoder named BioVIL-T [6], which is designed for LRRG task specifically. It is important to note that BioVIL-T utilizes image inputs of 448 x 448 pixels. Our method and other mainstream approaches use a standard input size of 224 x 224 pixels. The larger the image resolution, the richer the image details captured, which naturally provides an additional advantage to BioVIL-T. As illustrated in Fig. 7 (b), despite BioVIL-T’s use of larger images, our method still surpasses BioVIL-T. Additionally, as shown in Fig. 7 (b), the vanilla methods underperform in generating multi-period reports. In contrast, our method not only enhances accuracy in single-period report generation but also significantly improves multi-period report generation.
@@
Comparison with medical large language models. Currently, the field of large language models is rapidly evolving, leading to the emergence of numerous substantial medical models aimed at solving a variety of medical image-text tasks. To more comprehensively assess the value of our method, we conducted a comparison with existing medical large language model approaches, including RadFM [49], LLaVA-Med [50], Med-Flamingo [51] and MedDr [52]. The evaluation metrics follow those used in MedDr. As indicated in Tab. IV, in the RRG task, our specialized model demonstrates a distinct advantage over these larger models, particularly in terms of the RadCliQ score [63] (a comprehensive metric that integrates multiple metrics). This suggests that while large models possess robust generalizability and the capability to handle multiple medical image-text tasks, their performance in specific tasks is still inferior to that of dedicated models. Additionally, large models need longer training and inference times, as well as more hardware requirements. In practical applications, particularly for RRG tasks, large models currently cannot fully replace task-specific models.
IV-C Ablation Study and Case Study
Effect of multi-stage fusion. We initially validated the method utilizing multi-stage fusion. Our analysis compared two approaches: one that aligns multimodal features and detects differences only at the final visual stage, and another that applies feature alignment and difference detection across all visual stages. This comparison helps highlight the difference between our early fusion strategy that integrated within the visual encoder and the late fusion approach adopted by previous LRRG methods. As shown in Tab. III, compared to single-stage feature integration, multi-stage fusion achieved an increase in both the NLG and CE metrics (0.033$\uparrow$ on F1 of MIMIC-CXR and 0.012$\uparrow$ on F1 of Longitudinal-MIMIC), highlighting the importance of leveraging multi-level visual features for enhancing longitudinal representation learning.

### v2

@@ Abstract
Radiology Report Generation (RRG) automates the creation of radiology reports from medical imaging, enhancing the efficiency of the reporting process. Longitudinal Radiology Report Generation (LRRG) extends RRG by incorporating the ability to compare current and prior exams, facilitating the tracking of temporal changes in clinical findings. Existing LRRG approaches only extract features from prior and current images using a visual pre-trained encoder, which are then concatenated to generate the final report. However, these methods struggle to effectively capture both spatial and temporal correlations during the feature extraction process. Consequently, the extracted features inadequately capture the information of difference across exams and thus underrepresent the expected progressions, leading to sub-optimal performance in LRRG. To address this, we develop a novel dynamic difference-aware temporal residual network (DDaTR). In DDaTR, we introduce two modules at each stage of the visual encoder to capture multi-level spatial correlations. The Dynamic Feature Alignment Module (DFAM) is designed to align prior features across modalities for the integrity of prior clinical information. Prompted by the enriched prior features, the dynamic difference-aware module (DDAM) captures favorable difference information by identifying relationships across exams. Furthermore, our DDaTR employs the dynamic residual network to unidirectionally transmit longitudinal information, effectively modeling temporal correlations. Extensive experiments demonstrated superior performance over existing methods on three benchmarks, proving its efficacy in both RRG and LRRG tasks. Our code is published at https://github.com/xmed-lab/DDaTR
{IEEEkeywords}
@@ 1 Introduction
  • We propose a novel dynamic difference-aware temporal residual network (DDaTR) to better encode longitudinal information by adaptively learning the multi-level abstractions of different prior features and modeling the unidirectional information flow.
@@
2 Related Work
2.1 Radiology Report Generation
@@
2.2 Longitudinal Radiology Report Generation
@@
Although adding longitudinal features could alleviate the hallucinations about unobserved priors when training language models for report generation, there are still limitations in capturing the important longitudinal correlations. Firstly, they overlook the different contributions of previous image content (e.g., different viewpoints, limited longitudinal information or varying time intervals between examinations). This limitation leads to insufficient global temporal correlations. Secondly, previous feature fusion is used after visual encoder. Specifically, they use general visual pre-training encoders (e.g.,ResNet [30], SwinTransformer [31]) to extract visual features of each images independently, and then they leverage cross-attention based modules or transformer blocks to encode longitudinal features. However, the encoding process does not consider the relationships between multi-level features of longitudinal images, especially the low-level features including important details and texture information. This limitation leads to inaccurate local temporal correlations. Thirdly, previous methods fail to account for the prior knowledge in previous report during the visual feature extraction. This report documents the patient’s clinical findings and diagnoses, reflecting the underlying clinical intent of the examination, which is important for perceiving prior visual semantics.
3 Proposed Method
@@
3.1 Problem Definition and Our Baseline Framework
@@
$\small F_{img}=E_{img}\left(I_{n}\right)\vskip-5.69054pt$ (1)
@@
$F_{img}=\begin{cases}E_{ddatr}\left(I_{n}\right)&\text{if }n=1\\ E_{ddatr}\left(I_{n},I_{n-1},R_{n-1}\right),&\text{if }n>1\end{cases}\vskip-2.84526pt$ (4)
@@
3.2 Overview of Our Proposed DDaTR
@@
where $\hat{F}_{prior}^{m-1}$ represents the final output feature from the m1m-1 stage and $E_{prior}^{m}$ denotes the $m^{th}$ stage of prior visual encoder. For each stage of the prior visual encoder, we introduce the Dynamic Feature Alignment Module (DFAM) to dynamically fuse the image feature $F_{prior}^{m}$ and text feature $F_{txt}$, resulting in the aligned output feature $F_{align}^{m}$.
@@
where the mechanics of the $DFAM(\cdot)$ are detailed in Sec. 3.3. Next, this aligned output is added to $F_{prior}^{m}$ via a residual connection to obtain the final prior visual feature. The formula is given by:
@@
The mechanics of the $DDAM(\cdot)$ are detailed in Sec. 3.4. This output is also added to $F_{cur}^{m}$ using a residual connection, subsequently feeding into the next stage. Ultimately, this process culminates in the acquisition of the current visual feature $\hat{F}_{cur}^{m}$. The equation can be expressed as:
@@
Table 1: Comparison with existing report generation methods on MIMIC-CXR, Longitudinal-MIMIC and IU-XRay. * indicates results evaluated by us. For the MIMIC-CXR and IU-Xray, we adopt the evaluation outlined by RECAP[8], using the Macro score for CE metrics. For the Longitudinal-MIMIC, we follow L-R2Gen [3], calculating the Micro score for CE metrics.
@@
3.3 Dynamic Feature Alignment Module
@@
3.4 Dynamic Difference-aware Module
@@
Secondly, the dynamic fusion part of the DDAM is the same as DFAM, designed to ensure that the difference-aware features do not overwhelm the current visual features while allowing a controlled flow of longitudinal information into the next stage. This module dynamically integrates prior information, allowing the model to adaptively select the most relevant features. Specifically, $F_{cur}^{m}$ represents the visual features, $F_{da}$ means the difference-aware visual features including longitudinal features. When we fuse the $F_{cur}$ and $F_{da}$, we dynamically add $F_{da}$ to $F_{cur}$ using a gating mechanism. $f_{dy}$ represents our gating module, and the fusion output is $F_{dy}=f_{dy}(F_{da})$, where $F_{dy}$ represents the output weights generated by our gating module. Then, using $\alpha$’ indicates whether the input has a prior examination; it is set to 1 if a prior exists, and 0 otherwise. So the fused process is as follows:
@@
4 Experiments
4.1 Datasets and Metrics
@@
Experiment Settings. Our code, based on the Pytorch framework [47], was trained using a single NVIDIA RTX A40 GPU. For all methods used in our comparison, we strictly followed the default configurations reported in their respective original papers without any modifications. To ensure fair and consistent evaluation, we adopted the same training settings as the baseline PromptMRG across all our experiments, including those conducted on different datasets and with various visual backbones. Specifically, all input images were resized to 256 × 256, followed by data augmentations such as 224 × 224 random cropping and random rotation within ±5 degrees. The maximum input text length was set to 100 words. We employed the AdamW optimizer [48] with a weight decay of 0.05. The initial learning rate was set to 5e-5, and all models were trained for 10 epochs with a batch size of 16. Overall loss function combines the Language Modeling (LM) loss $\mathcal{L}_{LM}$ [49] for generation and standard Cross-Entropy (CE) loss $\mathcal{L}_{CE}$ for current and prior disease classification. The total loss $\mathcal{L}$ is defined as follows:
$\mathcal{L}=\mathcal{L}_{LM}+w(\mathcal{L}_{CE}^{cur}+\mathcal{L}_{CE}^{prior}),$ (21)
where $w$ is set to 4, consistent with the setting used in PromptMRG.
Table 2: Performance Comparison on three different visual backbones: ResNet101, SwinT-B, and EfficientNetv2. Vanilla represents that visual features were extracted using original pretrained models, coupled with the frozen pretrained BERT for text feature extraction, with subsequent concatenation of these features.
Datasets
Visual
Backbone
Methods Clinical Efficacy Natural Language Generation
@@
MIMIC
-CXR
ResNet101 Vanilla 0.395 0.426 0.389 0.108 0.156 0.267
@@
Longitudinal
-MIMIC
ResNet101 Vanilla 0.487 0.544 0.486 0.101 0.151 0.261
@@
Table 3: Ablation Study. w/o MSF indicates without multi-stage fusion. w/o DAM indicates without difference-aware module. w/o FAM indicates without feature alignment module. w/o DF indicates without dynamic fusion of both DFAM and DDAM.
@@
4.2 Experimental Results
Comparison on both RRG (MIMIC-CXR, IU-XRay) and LRRG (Longitudinal-MIMIC) benchmarks. We compare our method with a wide range of classical and SOTA methods on MIMIC-CXR, IU-XRay and Longitudinal-MIMIC, and report their results in Tab. 1. For the MIMIC-CXR dataset, we follow the evaluation established by RECAP, calculating the Macro score for the CE metrics. For the IU-Xray dataset, we use the same setting as PromptMRG that did not conduct any training or fine-tuning on it, but directly tested our model which trained on MIMIC-CXR. For the Longitudinal-MIMIC dataset, our approach follows the evaluation methodology used in LR2Gen, where the Micro score is calculated for the CE metrics. By comparison, our method achieves SOTA on CE metrics, significantly outperforming other approaches on both RRG and LRRG. We performed a statistical significance analysis on the MIMIC-CXR test set by calculating the p-value of the F1 scores across all longitudinal cases. The result is $P_{F1}$=0.0415 <0.05 comparing with previous SOTA promptmrg, indicating that the performance improvement of our method is statistically significant. This not only demonstrates that utilizing longitudinal information for report generation tasks yields reports that more closely align with the groundtruth, but also highlights our method’s capability to more accurately capture the diagnostic intent of the radiologists. For example, our method with SwinT-B on MIMIC-CXR achieves 0.441 of F1_macro CE score with an absolute improvement of 0.048. Our method with SwinT-B on Longitudinal-MIMIC achieves 0.527 of F1_micro CE score with an absolute improvement of 0.035.
Table 4: Performance Metrics on MIMIC-CXR Dataset. Comparison with Large language models.
@@ 4.2 Experimental Results
Table 5: Performance on ReXrank Dataset. Comparison with existing specialist RRG Models.
@@
Figure 7: (a) Illustrates different test settings. Each column represents a specific test setting, with n% indicating the proportion of longitudinal reports used. The total number of samples for each setting is shown at the top of each column. (b) Performance Comparison on different test settings.
@@
We analyze this as follows: a review of existing methods reveals that higher NLG scores often result from enhancements in the decoder. Thus, under similar data processing and training strategies, the NLG scores largely depend on the ability of the model’s decoder, which includes aspects such as the architecture, the tokenizer, and the pre-trained models. Our research primarily focuses on how to encode longitudinal information effectively, aligning input features more closely with medical intent. Consequently, we did not make specific structural modifications to the decoder but instead followed the decoder used in PromptMRG. Furthermore, our zero-shot results on IU X-Ray also outperform previous SOTA, which does not contain longitudinal information. This highlights the robustness and generalizability of our approach and further demonstrates that our model not only substantially improve longitudinal report generation, but also maintain the advanced performance in single-report scenarios.
@@
Comparison on different visual backbones. Our proposed algorithmic framework is adaptable to all multi-scale visual feature extraction networks. To validate its effectiveness, we conducted comparative experiments using three prominent visual backbones: ResNet101 [30], SwinT-B [31], and EfficientNetv2 [60]. These networks were selected because they are commonly employed for visual feature extraction in prior RRG research. We compared the performance outcomes between using standard vanilla visual networks and our method under the same baseline framework. In the vanilla approach, visual features were extracted using the original network architectures (e.g., ResNet101) and also fine-tuned during training using the same strategy as ours. For textual features, a pretrained BERT model was also used in a frozen state, and the visual and textual features were subsequently concatenated for downstream tasks. Experimental results, presented in Tab. 2, demonstrate that our method outperforms the vanilla encoder across all three backbones by a significant margin.
@@
Comparison on ReXrank. To compare our method with the latest approaches in RRG, we evaluated DDaTR on the official test set of MIMIC-CXR provided by the ReXrank leaderboard [61], which including the comprehensive metrics, such as BERTScore[62], SembScore[32], and GREEN[63]. The results are presented in Tab. 5. Our model ranks second on the leaderboard. While MedVersa [59] significantly outperforms all other methods on the leaderboard, it is trained on 10 public datasets and built upon LLM. Due to the substantial differences in both data amount and model size, direct comparison is not entirely fair. Compared to other task-specific LLMs that are explicitly trained on RRG data—such as LLM-CXR[56], CheXagent[58], RaDialog[55] and MAIRA-2[57]—our method achieves consistently superior performance across all evaluation metrics, further demonstrating the effectiveness of our approach under a more lightweight setting.
@@
Comparison on complexity. To compare model complexity, we report the trainable parameters for ours and PromptMRG as follows: PromptMRG (220M), Ours with Swin-T (360M), Ours with EfficientNetV2 (247M), and Ours with ResNet101 (419M). Notably, when using EfficientNetV2, our model maintains a comparable parameter count (247M) to PromptMRG (220M), yet still achieves significantly better performance as shown in Tab. 2. This indicates that the performance gains of our method are not merely attributed to increased model complexity. In fact, in Tab. 2 Longitudinal-MIMIC, ours with ResNet101 has the highest number of parameters but does not perform as well as the other backbones, further emphasizing that simply increasing model complexity does not guarantee better longitudinal modeling performance. Furthermore, PromptMRG is designed for single-period input, while our method handles longitudinal data, which is acceptable to introduce some additional parameters.
@@
Comparison on different proportions of longitudinal cases. We conducted assessments not only on standard MIMIC-CXR benchmark but also under various settings with different proportions of longitudinal cases. As shown in Fig. 7 (a), we split the testset of MIMIC-CXR. 0% indicates that the test data without longitudinal information. Conversely, 100% means all test cases. At 50%, the test set is composed by randomly selecting half of the reports with prior exams. In the above three testing settings, we evaluated the performance of different encoding methods within the baseline frameworks. Specifically, we tested the effectiveness of three prominent visual backbones using both the standard vanilla approach and our method, alongside another pretrained encoder named BioVIL-T [6], which is designed for LRRG specifically. It is important to note that BioVIL-T utilizes image inputs of 448 x 448 pixels. Our method and other mainstream approaches use a standard input size of 224 x 224 pixels. The larger the image resolution, the richer the image details captured, which naturally provides an additional advantage to BioVIL-T. As illustrated in Fig. 7 (b), despite BioVIL-T’s use of larger images, our method still surpasses BioVIL-T. Additionally, as shown in Fig. 7 (b), the vanilla methods underperform in generating multi-period reports. In contrast, our method not only enhances accuracy in single-period report generation but also significantly improves multi-period report generation.
@@
Comparison with medical large language models. Currently, the field of large language models is rapidly evolving, leading to the emergence of numerous substantial medical models aimed at solving a variety of medical image-text tasks. To more comprehensively assess the value of our method, we conducted a comparison with existing medical large language model approaches, including RadFM [50], LLaVA-Med [51], Med-Flamingo [52] and MedDr [53]. The evaluation metrics follow those used in MedDr. As indicated in Tab. 4, in the RRG task, our specialized model demonstrates a distinct advantage over these larger models, particularly in terms of the RadCliQ score [64] (a comprehensive metric that integrates multiple metrics). This suggests that while large models possess robust generalizability and the capability to handle multiple medical image-text tasks, their performance in specific tasks is still inferior to that of dedicated models. Additionally, large models need longer training and inference times, as well as more hardware requirements. In practical applications, particularly for RRG tasks, large models currently cannot fully replace task-specific models.
4.3 Ablation Study and Case Study
[…]